g. Penny’s age is one third her father’s age and in 13 years time her age will be half of her father’s age. How old is Penny now?
Answer by stanbon(75887)

Penny’s age is one third her father’s age and in 13 years time her age will be half of her father’s age. How old is Penny now?
----------
Equations:
f = 3p
2(p+13) = f+13
-----------------------
Substitute for "f" and solve for "p":
2p + 26 = 3p + 13
p = 13 (Penny's age now)
=============================
Cheers,
Stan H.
-----------------------------